An electric dipole, consisting of two opposite charges of 2 × 10-6 C each separated by a distance 3 cm is placed in an electric field of 2 × 105 N/C. Torque acting on the dipole is


(a)    12 × 10-1N - m


(b)    12 × 10-2N - m


(c)    12 × 10-3N - m


(d)    12 × 10-4N - m


Grade:11

4 Answers

Amit Saxena
10 years ago

(c)

Charges (q) = 2 × 10-6 C, Distance (d)

= 3 cm = 3 × 10-2 m and electric field (E)

= 2 × 105 N/C. Torque (t) = q.d.

E = (2 × 10-6) × (3 × 10-2) × (2 × 105)

= 12 × 10-3N-m.

Ajeet Tiwari
3 years ago
Hii Simran
The correcet answer is (c)
Charges (q) = 2 × 10^-6 C,
Distance (d)= 3 cm = 3 × 10^-2 m and
electric field (E)= 2 × 10^5 N/C.
Torque (t) = q.d.
E = (2 × 10^-6) × (3 × 10^-2) × (2 × 10^5)
= 12 × 10^-3N-m.
Thanks
Yash Chourasiya
3 years ago
Dear Student

Given,
Charges (q) = 2 × 10-6 C
Distance (d)= 3 cm = 3 × 10-2 m
and electric field (E)= 2 × 105 N/C.
We know that
Torque (t) = q.d. E
= (2 × 10-6) × (3 × 10-2) × (2 × 105)
= 12 × 10-3N-m.

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya
Ayush Sengar
2 years ago
Charge(q) = 2x10^-⁶ Coulomb, Distance(d) = 3 cm = 3x10^-